OpenStudy (wintersuntime):

Use what you know to about slope and y-intercept to graph the linear equations on the same coordinate plane provided at right.

OpenStudy (wintersuntime):

A. y=1/2x-8 m= b=

OpenStudy (wintersuntime):

B. y=1/2x+3 m= b=

OpenStudy (wintersuntime):

C. y=1/2x-3 m= b=

OpenStudy (wintersuntime):

D. y=1/2x m= b=

OpenStudy (wintersuntime):



OpenStudy (mathmale):

I'd suggest you enclose the slope, m, 1/2, in parentheses for clarify. D. y=(1/2)x What is the slope of D?

OpenStudy (wintersuntime):

1/2

OpenStudy (mathmale):

No calculations necessary; merely read the slope right off the equation. Right. m=1/2. What about the slopes of the other 3?

OpenStudy (kittiwitti1):

\[\text{slope-intercept format is }y=mx+b\]\[\text{slope is in format }\frac{\Delta y}{\Delta x}\text{, the y intercept is the variable }b\]

OpenStudy (wintersuntime):

For the slope of a its 1/2

OpenStudy (mathmale):

Yes. What about b? what about c?

OpenStudy (wintersuntime):

the same 1/2

OpenStudy (wintersuntime):

and for their y-intercept for a its 8 and for b its 3 and for c its 3

OpenStudy (wintersuntime):

For D is the y-intercept x ??

OpenStudy (wintersuntime):

so 0 or x?

OpenStudy (mathmale):

Can't say that D has no y-intercept. It does have one. Perhaps if you grph D, you'll see what the y-intercept is. Kittiwitti1: Nice work, only you are violating the Code of Conduct by doing others' work for them and handing out answers.

OpenStudy (kittiwitti1):

The second part was, yes. That's why I deleted it. I don't know why the first part was deleted though. It was just reformatting the problems.

OpenStudy (wintersuntime):

So for D is 0

OpenStudy (kittiwitti1):

@wintersuntime did you try graphing it out ?

OpenStudy (mathmale):

winter: "no y-intercept" actually points to a y-intercept of zero (0). What about the other 3? You had the numbers right, but there were some sign errors. Yes, b is 0 for D. Try again on a, b and c. look carefully at the original equations given.

OpenStudy (wintersuntime):

No

OpenStudy (mathmale):

D: the slope is ... ? and the y-intercept is b=?

OpenStudy (wintersuntime):

1.B=8 and 2. b= 3 3. b=3

OpenStudy (mathmale):

B=8? please explain.

OpenStudy (kittiwitti1):

Again, the slope format is \[y=mx+b\]If the y-intercept is -c we would get:\[y=mx+(-c)\]

OpenStudy (wintersuntime):

For letter b the slope is 3

OpenStudy (mathmale):

Aren't you supposed to find the slope, m, and the y-intercept, b (not B)?

OpenStudy (wintersuntime):

both

OpenStudy (mathmale):

I'm being picky, I know, but I want you to find the slope and y-intercept of B. Don't write B=8; that means nothing. You meant b=8.

OpenStudy (wintersuntime):

So am I suppose to graph all of them

OpenStudy (mathmale):

not necessry, but it never hurts to draw them for the practice. What are your answers to C? m= b=

OpenStudy (wintersuntime):

m=1/2 b=3

OpenStudy (mathmale):

Review your own answers. Then rethink C. What is m for C? What is b for C? Are you sure?

OpenStudy (wintersuntime):

-3

OpenStudy (mathmale):

The y-intercept of C is not b=3. Now you've got it; the y-int. for C is b=-3. Yes.

OpenStudy (wintersuntime):

oh i forgot to put the negative sign

OpenStudy (kittiwitti1):

Yes, make sure you put signs where necessary. Sometimes they're the deciding factor in the right or wrong answer (:

OpenStudy (mathmale):

Kitty: Is it necessary for you to inform us about your planned reinstatement? I think not. Your inputs are welcome so long as you guide the other person's work and do not hand out answers (or spam).

OpenStudy (mathmale):

Winter: Nice work. Done. Any questions?

OpenStudy (mathmale):

Yes, Ma'am.

OpenStudy (wintersuntime):

For graphing can I use the website called decimals

OpenStudy (mathmale):

Kitti: Your Witti might be Witti-er in the Math Chat, not in a serious discussion.

OpenStudy (mathmale):

Anything that correctly graphs these functions is fine. But I don't see the advantage of using decimals here.

OpenStudy (kittiwitti1):

I haven't commented in the math chat for several months, and I barely participate if at all... so I really have no idea what you are talking about.

OpenStudy (wintersuntime):

The website is called decimals were u could graph

OpenStudy (kittiwitti1):

ANYWAY back to the topic, sorry I got off track. Are you referring to the website desmos?

OpenStudy (mathmale):

I'm not famiar with that site, Winter. Try it and evaluate its worth.

OpenStudy (wintersuntime):

yes lol

OpenStudy (mathmale):

desmos seems to be popular with at least some openstudy users.

OpenStudy (wintersuntime):

cuz I'm bad at graphing

OpenStudy (kittiwitti1):

That's alright, go ahead and use it. (:

OpenStudy (wintersuntime):

I have a last question

OpenStudy (mathmale):

Thanks for your participation. Time for me to get off. Please actually try desmos and/or other graphing utilities and evaluate their worth to y ou. Last question? Go ahead.

OpenStudy (wintersuntime):

What would be the equation of a line parallel to the lines in parts (a) - (d) with a y-intercept of (0,100) ?

OpenStudy (mathmale):

Let me turn around and ask you this: What would the y-intercept be? b=?

OpenStudy (wintersuntime):

100

OpenStudy (wintersuntime):

y=0x+100?

OpenStudy (kittiwitti1):

i) parallel lines have equal slopes ii) the y intercept is 100, as you've said

OpenStudy (mathmale):

Right. You'd have a line parallel to the other four lines but with a y-intercept of 100.

OpenStudy (mathmale):

Au revoir. Auf wiedersehen.

OpenStudy (wintersuntime):

french

OpenStudy (kittiwitti1):

Ciao, MM. Have a good day o/

OpenStudy (wintersuntime):

धन्यवाद dhanyavaad

OpenStudy (wintersuntime):

thanks so much for ur helps guys

OpenStudy (kittiwitti1):

That's actually French and German ☺ But, anyway. You would have equal slope for the line parallel, and the point given, (0,100) is the y-intercept. That should be enough information.

OpenStudy (kittiwitti1):

No problem, good luck Winter! ♣

OpenStudy (wintersuntime):

thank you so much for u help once again

OpenStudy (kittiwitti1):

^-^